# CantonMT: Cantonese to English NMT Platform with Fine-Tuned Models using Synthetic Back-Translation Data

### v1

@@ Abstract
Neural Machine Translation (NMT) for low-resource languages is still a challenging task in front of NLP researchers. In this work, we deploy a standard data augmentation methodology by back-translation to a new language translation direction Cantonese-to-English. We present the models we fine-tuned using the limited amount of real data and the synthetic data we generated using back-translation including OpusMT, NLLB, and mBART. We carried out automatic evaluation using a range of different metrics including lexical-based and embedding-based. Furthermore. we create a user-friendly interface for the models we included in this CantonMT research project and make it available to facilitate Cantonese-to-English MT research. Researchers can add more models into this platform via our open-source CantonMT toolkit https://github.com/kenrickkung/CantoneseTranslation.
CantonMT: Cantonese to English NMT Platform with Fine-Tuned Models using Synthetic Back-Translation Data
@@ 1 Introduction
Cantonese is one of the most popular dialects of Chinese languages, after the standard language Mandarin (the current official language in China, originally from the Beijing area), originally from the capital of Guangdong province, Guangzhou (a.k.a. Canton) in China. The population of Guangdong province is 126.84 million in 2021. In addition, Cantonese is also the native language in Hong Kong and Macau regions which have populations of 7,503,100 and 704,149 in 2023, according to HK Census and Statistics Department https://www.censtatd.gov.hk/en/ and Macrotrends Global Population statistics https://www.macrotrends.net/global-metrics/countries/MAC/macao. Furthermore, because of the economic growth in Guangdong, HK, and Macau, many people from other Chinese provinces also learned to speak Cantonese for job purposes and cultural influences. There is also a big amount of population worldwide outside of China speaking Cantonese. In the era of the fast development of natural language processing (NLP), many machine translation (MT) models have been proposed for the majority of languages worldwide. However, low-resource language MT still challenges researchers. Cantonese translation using MT, among one of them, is yet to be developed without much attention being paid so far.
In this work, we aim at investigating one of the popular methods, i.e. synthetic data augmentation via back-translation and model fine-tuning, on Cantonese-to-English neural MT (NMT), a new language pair. We select several models for comparisons including both smaller and larger language models. We compare the system performance using a range of evaluation metrics. Furthermore, we open source our toolkit and create a web-based user-friendly platform called CantonMT to facilitate the research on Cantonese-English translation and beyond. 111a video demo of CantonMT is available at https://youtu.be/s8P5fJjS7Ls
In the next section (Section 2), we survey related work on Cantonese-English MT, data augmentation for MT, and available demos/engines. Section 3 introduces our methodology and framework. Section 4 explains the web-based CantonMT platform. Section 5 concludes this work with discussions.
@@ 2 Related Work
Research work focusing on Cantonese-English MT has not gained much attention to date. Earliest works include Wu et al. (2006) where Example-based and Rule-based MT were investigated. In recent years, there was the project plan on Cantonese-English Translation at HKU by Wing (2020) where the authors listed models to be investigated including RBMT, EBMT, SMT, GRU, Transformer models. In a looser connection, regardless of English as the target language, there has been some work on Cantonese-related MT. These include dialectal translation between Cantonese and Mandarin Chinese by Zhang (1998), Yi Mak and Lee (2022), and Liu (2022).
Data augmentation via Backtranslation has been one of the standard practices to generate a synthetic corpus for assisting the MT performances of low-resource language pairs. This has been popular for both SMT and NMT Sugiyama and Yoshinaga (2019); Graça et al. (2019); Edunov et al. (2020); Nguyen et al. (2021); Pham et al. (2023). However, to the best of our knowledge, none of these works focused on Cantonese-to-English translation.
Existing platforms or off-the-shelf demos for Cantonese-to-English MT are very Scarce. Popular MT engines from commercial IT companies do not include this language pair including Google Translator https://translate.google.com and Bing Translator https://www.bing.com/translator. Both of them only included simplified and traditional characters of Mandarin Chinese. The available translators from commercial IT companies include Baidu Translator (Fanyi) https://fanyi.baidu.com/, an IT company from China that includes several Chinese dialectal languages  [^2] All these websites were last visited 2024 March 4th. In the opposite direction, there are open-source tools on English-to-Cantonese MT from TransCan https://github.com/ayaka14732/TransCan.
@@ 3 Experimental Work
We introduce the methodology of CantonMT, experimental evaluations using 38K words.hk real bilingual corpus, and extended work when we acquired more real bilingual dictionary data from Wenlin.com.
@@ 3.1 Methodology and Framework
  • ModelFineTunePhase2: second step MT fine-tuning using real and synthetic data (ft-syn, v2)
@@
For data collection, we scraped the data from the public Hong Kong forum LIHKG https://lihkg.com, which was launched in 2016 and has multiple categories including sports, entertainment, hot topic, gossip, current affairs, etc. We extracted more than 1 million sentences from this website, however, the raw data is with a lot of noises that need to be cleaned, with an example shown in Figure 6 (Appendix Section A). We did data cleaning to reduce noisy strings as well as data anonymisation by removing the user IDs from the text. We also filtered out the sentences that were too short with less than 10 Chinese characters. In the end, we prepared 200K clean monolingual Cantonese sentences for parallel synthetic data generation purposes. We shuffled the data for model training.
@@
In the model fine-tuning phase 1, we aim to train a set of reasonable Cantonese-English MT models for synthetic data generation and model comparisons. The baseline models we selected are Opus-MT, NLLB, and mBART with the following rationale: 1) How much does the model size impact the fine-tuning performances? For this, we use Opus-MT which is a much smaller model trained on Opus corpus using MarianMT framework and NLLB-200, a very large language model pre-trained on 200+ languages from Mata-AI. 2) How much does it matter if the pre-trained translation models have Cantonese in their pre-training? For this, we add mBART (mbart-large-50-many-to-many-mmt) which is another LLM but without Cantonese in his pre-training, vs NLLB which has Cantonese. Because the full-size NLLB is too large, we used the distilled model “nllb-200-distilled-600M”.
@@
We fine-tuned these models using the available bilingual data from a bilingual Cantonese-English dictionary “Yue-Dian” from https://words.hk, which is in total 44K in size. We divided this data into (training, development, and testing) sets with the sizes (38K, 3K, and 3K), in light that the standard WMT –workshop of machine translation – shared task uses around 3K sentences for testing sets.
@@
In Step 3 for synthetic data generation, we used Step 2 initially fine-tuned LLMs (LLM-ft-v1) to translate the collected monolingual Cantonese text we extracted from Step 1. In this way, we get 200K back-translated English sentences and these synthetic sentences together with the Cantonese sentences create the 200K synthetic parallel corpus we generated. From now on, we call the synthetic parallel corpus 200K-ParaSyn.
@@
In Step 4, we apply different ratios on the real parallel data we have and the 200K-ParaSyn for LLM fine-tuning. We also test the influence of model switches, i.e. using different types of LLMs for LLM-ft (Phase 1) and LLM-syn (Phase 2).
@@
In the last step, we deploy the fine-tuned LLMs in Phase 2 (LLM-syn) on the same test data and compare the results with LLM-ft (Phase 1) and baseline models without fine-tuning. We also report comparisons with available translation engines from IT companies, such as Baidu Translator, Bing Translator, and GPT4. For GPT4, there is an available fine-tuned version towards translation to Cantonese by “community builder”  [^3] https://chat.openai.com/share/7ee588af-dc48-4406-95f4-0471e1fb70a8 called Cantonese Companion.
@@
We used a range of different evaluation metrics including lexical-based SacreBLEU Post (2018), hLEPOR Han et al. (2013a, 2021), and embedding-based BERTscore Zhang* et al. (2020) and COMET Rei et al. (2020). hLEPOR has reported much higher correlation scores to the human evaluation than BLEU and other lexical-based metrics on the WMT shared task data Han et al. (2013b). However, recent WMT metrics task findings have demonstrated the advantages of neural metrics based on embedding space similarities Freitag et al. (2022).
@@
Figure 2: Learning Curve of Model Training using Real Data
@@
Figure 3: Automatic Evaluation Scores from Different Models in CantonMT. bl: bilingual real data; syn: synthetic data; h:h - half and half; 1:1/3/5 - 100% real + 100/300/500% synthetic; 10E: 10 epochs (default: 3); top-down second slot: model switch: model type using NLLB but synthetic data from other models (mBART and OpusMT); top-down third slot: including model switch for mBART fine-tuning using synthetic data generated from NLLB; similarly top-down forth slot: including model switch for OpusMT fine-tuning using synthetic data from NLLB. Bottom slot of Cluster 1: Bing/Baidu Translator and GPT4-finetuned Cantonese Companion; bold case is the best score of the same slot among the same model categories. Cluster 2: bilingual fine-tuned models using 38K words.hk data plus 14.5k Wenlin data; italic indicates the number outperforms the same model fine-tuned with less data 38K.
@@ 3.2 Evaluations of CantonMT
The Learning curves of three base models during training using the 38K real data are shown in Figures 2 from left to right: mBART, NLLB-200, and Opus-MT. We used three epochs for mBART because it was too large for our available computational resources. From the learning curves, we can see that NLLB-200 has a peak score at Epoch 3 then there is a dramatic drop till Epoch 6 and grows back at Epoch 10. Instead, the Opus-MT model has a steady increasing of the SacreBLEU score with more epochs though there are litter drops in between.
The automatic evaluation scores from CantonMT models and other commercial engines are listed in Figure 3. There are some interesting findings from the evaluation outcomes.
  • LLM-ft vs -LLM-ft-syn: 1) NLLB-syn-1:1 has slightly better scores than NLLB-bl on all metrics, but more ratio of synthetic data will decrease the scores such as 1to3 and 1to5 with around 1 absolute SacreBLEU point. 2) Similarly, mBART-syn-1:1 also outperforms mBART-ft but more ratios of synthetic data will reduce the evaluation scores such as 1to3. 3) Surprisingly, the synthetic model for Opus-mt does not win the Opus-ft-bl, which indicates the quality of the generated synthetic data matters.
  • Model Switch Matters: 1) the NLLB fine-tuned model using synthetic data from mBART (top-secon slot) produced higher scores than using the synthetic data generated from its own (top-first slot). 2) mBART fine-tuned using NLLB-generated synthetic data also outperforms mBART fine-tuning using only bilingual real data. 3) In a similar situation, Opus-MT perform differently in comparison to the other two models.
  • MT from IT companies: 1) GPT4-finetuned produced the highest evaluation scores but the free version of GPTs restricts the input number of strings; furthermore, it is much more unclear how GPT4 performed the MT; in addition, there are risks on data privacy when users choose to use engines from commercial companies. In contract, CantonMT is open-sourcely free and researchers can continue to fine-tune it with their data, or insert more models, fully confidential for users. 2) Bing and Baidu translators produced similar evaluation scores to the best system from CantonMT, though Bing produced slightly higher scores than Baidu, especially on lexical-based metrics SacreBLEU and hLEPOR.
  • Comparing to Model Deployment without Finetuning: in Cluster 3 (bottom) of the score table, model deployment without fine-tuning has much lower scores; these scores show that the fine-tuning and synthetic data augmentation make a large ratio of score increase of around 50% for all models using SecreBLEU.

### v3

# CantonMT: Cantonese to English NMT Platform with Fine-Tuned Models using Real and Synthetic Back-Translation Data

@@ Abstract
Neural Machine Translation (NMT) for low-resource languages remains a challenge for many NLP researchers. In this work, we deploy a standard data augmentation methodology by back-translation to a new language translation direction, i.e., Cantonese-to-English. We present the models we fine-tuned using the limited amount of real data and the synthetic data we generated using back-translation by three models: OpusMT, NLLB, and mBART. We carried out automatic evaluation using a range of different metrics including those that are lexical-based and embedding-based. Furthermore, we create a user-friendly interface for the models we included in this project, CantonMT, and make it available to facilitate Cantonese-to-English MT research. Researchers can add more models to this platform via our open-source CantonMT toolkit, available at https://github.com/kenrickkung/CantoneseTranslation.
CantonMT: Cantonese to English NMT Platform with Fine-Tuned Models using Real and Synthetic Back-Translation Data
@@ 1 Introduction
Cantonese is one of the most popular dialects of Chinese languages, after the standard language Mandarin (the current official language in China, originally from the Beijing area), originally from the capital of Guangdong province, Guangzhou (a.k.a. Canton) in China. The population of Guangdong province was 129.51 million in 2022 according to the National Bureau of Statistics of China  [^1] https://data.stats.gov.cn/english. In addition, Cantonese is also the native language in Hong Kong (HK) and Macau regions which have populations of 7,503,100 and 704,149 in 2023, according to HK Census and Statistics Department [^2] https://www.censtatd.gov.hk/en/ and Macrotrends Global Population statistics.333 https://www.macrotrends.net/global-metrics/countries/MAC/macao Furthermore, because of the economic growth in Guangdong, HK and Macau, many people from other Chinese provinces also learned to speak Cantonese for job purposes and due to cultural influences. There is also a large global population outside of China speaking Cantonese, 85.5 million, according to the Cantonese Language Association (CLA)  [^4] https://cantoneselanguageassociation.byu.edu/. In the era of the fast development of natural language processing (NLP), many machine translation (MT) models have been proposed for the majority of languages worldwide. However, low-resource language MT remains a challenge for researchers. Cantonese translation using MT, specifically, is under-explored and has not been given much attention thus far.
In this work, we investigate one of the more popular MT methods, i.e. synthetic data augmentation via back-translation and model fine-tuning, as an approach to Cantonese-to-English neural MT (NMT), along the way introducing Cantonese-English as a new language pair. We select several models for evaluation including both smaller and larger language models, and compare their system performance using a range of evaluation metrics. Furthermore, we open-source our toolkit and create a web-based user-friendly platform called CantonMT to facilitate research on Cantonese-English translation. A public video demo is available. [^5] CantonMT demo https://youtu.be/s8P5fJjS7Ls
In the next section (Section 2), we survey related work on Cantonese-English MT, data augmentation for MT, and available demos/engines. Section 3 introduces our methodology and framework. Section 4 explains the web-based CantonMT platform. Section 5 concludes this work with a discussion.
@@ 2 Related Work
Research work focussing on Cantonese-English MT has not gained much attention to date. Earliest efforts include the work of Wu et al. (2006) where example-based and rule-based MT were investigated. In recent years, a project plan on Cantonese-English Translation was put forward by researchers at the University of Hong Kong (HKU) where they proposed to investigate various MT approaches, including rule-based MT (RBMT), example-based MT (EBMT), statistical MT (SMT), gated-recurrent units (GRU) and transformers Wing (2020). More loosely related work include research in MT for Cantonese, but without English as the target language. These include dialectal translation between Cantonese and Mandarin Chinese by Zhang (1998), Yi Mak and Lee (2022) and Liu (2022).
Data augmentation via backtranslation has been one of the standard practices for generating a synthetic corpus for improving MT performance on low-resource language pairs. This has been popular for both statistical MT (SMT) and NMT Sugiyama and Yoshinaga (2019); Graça et al. (2019); Edunov et al. (2020); Nguyen et al. (2021); Pham et al. (2023). However, to the best of our knowledge, none of these efforts focused on Cantonese-to-English translation.
Existing platforms or off-the-shelf demos for Cantonese-to-English MT are very scarce. Popular MT engines from commercial IT companies, including Google Translate [^6] https://translate.google.com and DeepL Translator, [^7] https://www.deepl.com/translator do not include this language pair. Both of them only included simplified and traditional characters of Mandarin Chinese. Meanwhile, Microsoft Bing Translator [^8] https://www.bing.com/translator and Baidu, an IT company from China, made the Baidu Translator (Fanyi) [^9] https://fanyi.baidu.com/ available, which includes Cantonese among several Chinese dialectal languages. [^10] All these websites were last visited 4th March 2024. In the opposite direction, there are open-source tools for English-to-Cantonese MT from TransCan. [^11] https://github.com/ayaka14732/TransCan
@@ 3 Experimental Work
We introduce the methodology of CantonMT, experimental evaluations using the initial 38K real bilingual corpus, and extended model evaluations when we acquired 14.5K and 10K more real bilingual data from different sources subsequently.
@@ 3.1 Methodology and Framework
  4. 4.
    ModelFineTunePhase2: second step MT fine-tuning using real and synthetic data (ft-syn)
  5. 5.
@@
For data collection, we scraped the data from the public Hong Kong forum LIHKG, [^12] https://lihkg.com which was launched in 2016 and has multiple categories including sports, entertainment, hot topic, gossip, current affairs, etc. We extracted more than 1 million sentences from this website; however, the raw data comes with a lot of noise that needs to be cleaned, an example of which is shown in Figure 5 of Appendix A. We carried out data cleaning to reduce noisy strings as well as data anonymisation by removing user IDs from the text. We also filtered out the sentences that were too short, i.e., with less than 10 Chinese characters. In the end, we prepared 200K clean monolingual Cantonese sentences for parallel synthetic data generation purposes. We shuffled the data for model training.
@@
In model fine-tuning phase 1, we aim to train a set of reasonable Cantonese-English MT models for synthetic data generation and model comparisons. The baseline models we selected are Opus-MT, NLLB and mBART. These were chosen to answer the following questions: (1) How much does model size impact fine-tuning performance? For this, we use Opus-MT which is a much smaller model trained on the Opus corpus using the MarianMT framework and NLLB-200, a very large language model pre-trained on 200+ languages from Mata-AI; (2) To what extent does it matter if the pre-trained translation models are exposed to Cantonese in their pre-training? For this, we add mBART (mbart-large-50-many-to-many-mmt) which is another LLM but without Cantonese in its pre-training, vs NLLB which includes Cantonese. Because the full-size NLLB is too large, we used the distilled model nllb-200-distilled-600M.
@@
We fine-tuned these models using the available bilingual data from a bilingual Cantonese-English dictionary called “Yue-Dian”, [^13] https://words.hk which is in total 44K in size. We divided this data into training, development and testing sets with 38K, 3K and 3K as their respective sizes, in light of the fact that the shared tasks organised by the Workshop on Statistical Machine Translation (WMT) tend to include around 3K sentences in their test sets.
@@
In Step 3, synthetic data generation, we used the fine-tuned LLMs (LLM-ft-v1) from Step 2 to translate the monolingual Cantonese text we collected in Step 1. In this way, we obtain 200K back-translated English sentences; these synthetic sentences together with the Cantonese sentences create the 200K synthetic parallel corpus we generated. From now on, we will refer to the synthetic parallel corpus as 200K-ParaSyn.
@@
In Step 4, we apply different ratios on the real parallel data we have at hand and on 200K-ParaSyn for LLM fine-tuning. We also test the influence of model switches, i.e. using different types of LLMs for LLM-ft (Phase 1) and LLM-syn (Phase 2).
@@
In the last step, we deploy the fine-tuned LLMs in Phase 2 (LLM-syn) on the same test data and compare the results with LLM-ft (Phase 1) and baseline models without fine-tuning. We also report comparisons with commerically available translation engines such as the Baidu Translator, Bing Translator and GPT4. The implementation of GPT-4 that we used is Cantonese Companion, which was custom-made for translation to Cantonese by a community builder. [^14] https://chat.openai.com/share/7ee588af-dc48-4406-95f4-0471e1fb70a8
@@
We used a range of different evaluation metrics including the lexical-based SacreBLEU Post (2018) and hLEPOR Han et al. (2013a, 2021), and the embedding-based BERTscore Zhang* et al. (2020) and COMET Rei et al. (2020). hLEPOR has reported much higher correlation scores to the human evaluation than BLEU and other lexical-based metrics on the WMT shared task data Han et al. (2013b). However, recent WMT metrics task findings have demonstrated the advantages of neural metrics based on embedding space similarities Freitag et al. (2022).
(a) OPUS Baseline
(b) NLLB Baseline
(c) mBART Baseline
@@
Figure 2: Learning curves during model training using real data.
Model Name SacreBLEU hLEPOR BERTscore COMET
nllb-forward-bl 16.5117 0.5651 0.9248 0.7376
nllb-forward-syn-h:h 15.7751 0.5616 0.9235 0.7342
nllb-forward-syn-1:1 16.5901 0.5686 0.925 0.7409
nllb-forward-syn-1:1-10E 16.5203 0.5689 0.9247 0.738
nllb-forward-syn-1:3 15.9175 0.5626 0.924 0.7376
nllb-forward-syn-1:5 15.8074 0.562 0.9237 0.7386
nllb-forward-syn-1:1-mbart 16.8077 0.571 0.9256 0.7425
nllb-forward-syn-1:3-mbart 15.8621 0.5617 0.9246 0.7384
nllb-forward-syn-1:1-opus 16.5537 0.5704 0.9254 0.7416
nllb-forward-syn-1:3-opus 15.9348 0.5651 0.9242 0.7374
mbart-forward-bl 15.7513 0.5623 0.9227 0.7314
mbart-forward-syn-1:1-nllb 16.0358 0.5681 0.9241 0.738
mbart-forward-syn-1:3-nllb 15.326 0.5584 0.9225 0.7319
opus-forward-bl-10E 15.0602 0.5581 0.9219 0.7193
opus-forward-syn-1:1-10E-nllb 13.0623 0.5409 0.9164 0.6897
opus-forward-syn-1:3-10E-nllb 13.3666 0.5442 0.9167 0.6957
baidu 16.5669 0.5654 0.9243 0.7401
bing 17.1098 0.5735 0.9258 0.7474
gpt4-ft(CantoneseCompanion) 19.1622 0.5917 0.936 0.805
nllb-forward-bl-plus-wenlin14.5k 16.6662 0.5828 0.926 0.7496
mbart-forward-bl-plus-wenlin14.5k 15.2404 0.5734 0.9238 0.7411
opus-forward-bl-plus-wenlin14.5k 13.0172 0.5473 0.9157 0.6882
nllb-200-deploy-no-finetune 11.1827 0.4925 0.9129 0.6863
opus-deploy-no-finetune 10.4035 0.4773 0.9082 0.6584
mbart-deploy-no-finetune 8.3157 0.4387 0.9005 0.6273
nllb-forward-all3corpus 16.9986 0.583 0.927 0.7549
nllb-forward-all3corpus-10E 16.1749 0.5728 0.9254 0.7508
mbart-forward-all3corpus 16.3204 0.5766 0.9253 0.7482
opus-forward-all3corpus-10E 14.4699 0.5621 0.9191 0.7074
@@
Table 1: Automatic Evaluation Scores from Different Models in CantonMT. bl: bilingual real data; syn: synthetic data; h:h - half and half; 1:1/3/5 - 100% real + 100/300/500% synthetic; 10E: 10 epochs (default: 3); top-down second slot: model switch: model type using NLLB but synthetic data from other models (mBART and OpusMT); top-down third slot: including model switch for mBART fine-tuning using synthetic data generated from NLLB; similarly top-down forth slot: including model switch for OpusMT fine-tuning using synthetic data from NLLB. Bottom slot of Cluster 1: Bing/Baidu Translator and GPT4-finetuned Cantonese Companion; bold case is the best score of the same slot among the same model categories. Cluster 2: bilingual fine-tuned models using 38K words.hk data plus 14.5k Wenlin data; italic indicates the number outperforms the same model fine-tuned with less data 38K. Cluster 3: Deployed Model without fine-tuning Cluster 4: Finetuned with the previous 2 corpora and an additional 10K data from OPUS Corpora we managed to find in the end - it shows the evaluation improvement continues.
@@ 3.2 Evaluations of CantonMT
The learning curves of three base models during training using the 38K real data are shown in Figure 2 from left to right for mBART, NLLB-200 and Opus-MT. We used three epochs for mBART because it is too large for the computational resources available to us. From the learning curves, we can see that NLLB-200 has a peak score at epoch 3 then there is a dramatic drop until epoch 6, followed by an increase until epoch 10. In contrast, the Opus-MT model achieves a steady increase in its SacreBLEU score with more epochs, although there are little drops in between.
The automatic evaluation scores from CantonMT models and other commercial engines are listed in Table 3. Below are some interesting findings from the evaluation outcomes.
  • LLM-ft vs -LLM-ft-syn: (1) NLLB-syn-1:1 has slightly better scores than NLLB-bl on all metrics, but increasing the ratio of synthetic data will decrease the scores such as in the 1:3 and 1:5 configurations, with around 1 absolute SacreBLEU point. (2) Similarly, mBART-syn-1:1 also outperforms mBART-ft but increasing the ratio of synthetic data will reduce the evaluation scores such as in the 1:3 configuration. (3) Surprisingly, the synthetic model for Opus-mt does not outperform Opus-ft-bl, which indicates that the quality of the generated synthetic data matters.
  • Model Switching Matters: (1) the NLLB fine-tuned model using synthetic data from mBART (second model from the top of the table) produced higher scores than using the synthetic data generated from its own (first model from the top of the table). (2) mBART fine-tuned using NLLB-generated synthetic data also outperforms mBART fine-tuning using only bilingual real data. (3) In a similar situation, Opus-MT performs differently in comparison to the other two models.
  • Commercial MT models: (1) GPT4-finetuned produced the highest evaluation scores but the free version of GPTs restricts the input number of strings; the data size used for fine-tuning GPT-4 is unknown and such data is not publicly available to researchers; furthermore, it is unclear how GPT-4 performs MT; in addition, there are risks to data privacy when users choose to use engines from commercial companies. In contrast, CantonMT is open-source, free, and researchers can continue to fine-tune it with their data or include more models, and is fully confidential for users. 2) Bing and Baidu translators produced similar evaluation scores to the best system from CantonMT, though Bing produced slightly higher scores than Baidu, especially on the lexical-based metrics SacreBLEU and hLEPOR.
  • Comparing to Model Deployment without Finetuning: in Cluster 3 (bottom) of Table 3, model deployment without fine-tuning has much lower scores; these scores show that fine-tuning and synthetic data augmentation lead to a large increase in scores of around 50% for all models using SecreBLEU.